Oxley Machining Company has identified the following overhead activities, cost, and activity drivers for the coming year Expected Cost Acti Activity Capacity Activity Setting u Inspecting Grindin Receiving Number of setups 300 $45,000 Inspection hours 4,500 Machine hours 18,000 Number of parts 50,000 The company produces several different machine widgets used by other manufacturers. Information on separate batches for two of these widgets are: Widget A $850 $600 100 Widget B $950 $600 50 Direct materials Direct labour Units completed No. Qf setups Inspection hours Machine hours Parts used 20 20 30 40 The companys normal activity is 20,000 direct labour hours. Each batch uses 50 hours of direct labour. Upon investigation, you discover that Receiving employs one worker, who spends 80% of her time on the receiving activity and 20% of her time on inspecting products. Her salary is $40,000. Inspecting has taken into account the workers salary but not Receiving. In addition, Receiving also uses a forklift, at a cost of $8,000 per year for amortization and fuel. The forklift is used only in receiving.
Answer #1

13.A) The amount of cost assigned to receiving activity

=(40000*80%)+8000=40000

Correct option is C.

Since 80% of the employees time is spent on receiving activity and the forklift is only used for receiving, these have been included in computation.

13b) Activity rate for receiving

=40000/50000 =$ 0.8 per part

Correct option is A.

13 C)Activity rate for setting up equipment

=Setting up cost/number of set ups

=$60000/300

=$200 per set up

Correct option is E.

13D) The activity rate for grinding

=Grinding cost/machine hours

=$90000/18000

=$5 per machine hour

Correct option is C.

* First four sub parts have been answered as HOMEWORKLIB's policy.